# CLAPS: Posterior-Aware Conformal Intervals via Last-Layer Laplace

### v2

@@ 2 Related Work
Split conformal regression calibrates residual-shaped scores to guarantee marginal coverage, and fold-reuse variants such as jackknife+ and cross-fitting improve data efficiency (barber2021predictive; angelopoulos2023conformal). Recent work emphasizes efficiency—minimizing expected length or controlling a user-chosen risk—while preserving guarantees, including optimization-driven procedures and general risk-control frameworks (kiyani2024length; angelopoulos2022conformal; overman2024aligning). Concurrently, methods move beyond absolute residuals toward distribution-shaped scores that leverage estimates of P(Y|X)P(Y\,|\,X) (e.g., CDF-/quantile-based and probabilistic variants) or reframe regression as classification to better capture heteroscedasticity, skewness, and multimodality (chernozhukov2021distributional; romano2019conformalized; guha2024conformal; plassier2024probabilistic). We adopt the same rank-based split calibration rule but define the conformity via a symmetric CDF from a predictive distribution (instantiated later with LLLA), placing our approach on the distribution-shaped branch; for completeness we also compare to interval-thresholding baselines tailored for width efficiency (luo2025conformal).
@@
A major route to width efficiency under input-dependent noise is to learn scale or quantiles and calibrate them with conformal ranks. Conformalized Quantile Regression (CQR) replaces absolute-residual scores by a max-quantile deviation, achieving robustness to heteroscedasticity and skewed tails (romano2019conformalized). Building on this idea, boosted or reweighted schemes further reduce interval length by improving the base quantile fits while preserving split-calibrated guarantees (xie2024boosted). Localization and weighting adapt calibration to covariate neighborhoods or strata, yielding relaxed conditional guarantees and improved coverage balance across subpopulations (hore2023conformal; bhattacharyya2024group). When conditional distributions are multi-modal/asymmetric, reframing regression targets (e.g., via discretization or surrogate tasks) can enhance quantile learning before conformalization (guha2024conformal). Recent general treatments study heterogeneity and non-exchangeability directly, proposing efficient conformal procedures with validity under drift and data heterogeneity (plassier2024efficient). In survival/structured settings, conformalized quantile ideas are extended with monotonicity-aware or task-specific heads to stabilize learned quantiles while keeping rank calibration (qi2024toward). We adopt CQR and localized/weighted variants as primary baselines; our method differs by deriving a posterior-aware, CDF-symmetric score, whereas these approaches learn input-dependent scales/quantiles and then apply the same split rank rule.
@@
Posterior-aware conformal methods use a predictive distribution to define CDF/HPD-style conformity scores, then retain split/rank calibration. Recent advances show how modelling the conditional distribution improves conditional validity and efficiency, e.g., clustering- or distribution–driven calibration (zhang2024posterior; plassier2024probabilistic; chernozhukov2021distributional). In neural networks, a lightweight route to such predictive distributions is to endow only the final linear layer with a Bayesian treatment. Last-layer Laplace and its modern variants provide Gaussian posterior predictives around the MAP solution with negligible retraining cost (kristiadi2020being; daxberger2021laplace). Beyond the classical diagonal/empirical-Hessian forms, recent work improves the fidelity–cost trade-off via variational last layers, function-space priors, and geometry-aware approximations (harrison2024variational; cinquin2024fsp; bergamin2023riemannian). We adopt this last-layer Bayesianization to instantiate a symmetric CDF score from the LLLA posterior predictive on a shared MLP backbone, then calibrate by ranks; this places our method at the intersection of posterior-aware conformal scoring and efficient Laplace-based uncertainty, and is complementary to non-Bayesian distributional surrogates (zhang2024posterior; plassier2024probabilistic).
@@
Beyond score design, three complementary axes improve efficiency and coverage balance. First, localization tailors calibration to neighborhoods of the test covariate, yielding relaxed conditional guarantees and better subgroup balance (guan2023localized; hore2023conformal). Second, weighting focuses coverage where distribution shift or stratification matters—either by importance-style reweighting or group-aware corrections—without changing the base score (bhattacharyya2024group). Third, bootstrap ensembles aggregate diverse learners or resamples to stabilize interval length under model misspecification or dependence, including sequential/time-series settings (xu2021conformal; xu2023conformal; gupta2022nested; qian2024random; rivera2024conformal). These lines are orthogonal to our posterior-aware scoring: our method can plug into localized or weighted calibration, and can be paired with ensemble wrappers to trade off width and robustness while preserving split/rank guarantees.
@@ 3 Method
We propose CLAPS, a posterior-aware conformal regression method that combines a Last-Layer Laplace Approximation (LLLA) with split-conformal calibration. CLAPS aligns the nonconformity score with the model’s posterior predictive shape and then uses rank-based calibration to guarantee finite-sample coverage.
@@ 3.2 Last-Layer Laplace Approximation (LLLA)
We place a Gaussian prior $w\sim\mathcal{N}(0,\lambda^{-1}I)$ and assume homoscedastic Gaussian noise $y\mid x,w\sim\mathcal{N}(\phi(x)^{\top}w,\sigma^{2})$. Let $\Phi\in\mathbb{R}^{n\times d}$ be the design matrix of $\phi(x)$ for $\mathcal{D}_{\text{tr}}$ and $y\in\mathbb{R}^{n}$ the targets. The MAP solution and posterior covariance for the last layer are
$w_{\text{MAP}}$ $=\arg\min_{w}\tfrac{1}{2\sigma^{2}}\lVert y-\Phi w\rVert_{2}^{2}+\tfrac{\lambda}{2}\lVert w\rVert_{2}^{2},$
$\Sigma$ $=\bigl(\lambda I+\tfrac{1}{\sigma^{2}}\Phi^{\top}\Phi\bigr)^{-1}.$
The LLLA posterior is $w\mid\mathcal{D}_{\text{tr}}\sim\mathcal{N}(w_{\text{MAP}},\Sigma)$ and yields a Gaussian posterior predictive for any $x$,
$y\mid x,\mathcal{D}_{\text{tr}}$ $\sim\mathcal{N}\!\bigl(\hat{\mu}(x),\,\hat{v}(x)\bigr),$
$\hat{\mu}(x)$ $=\phi(x)^{\top}w_{\text{MAP}},$
$\hat{v}(x)$ $=\sigma^{2}+\phi(x)^{\top}\Sigma\,\phi(x).$
We fix the last–layer prior precision to $\lambda=1.0$ unless otherwise stated, and estimate the noise variance $\sigma^{2}$ either by maximizing the (approximate) marginal likelihood under the last–layer Gaussian model or—consistently in our experiments—via residuals on the training split $\mathcal{D}_{\mathrm{tr}}$ (with standardized target):
$\sigma^{2}$ $=\frac{1}{|\mathcal{D}_{\mathrm{tr}}|}\,\bigl\|y_{\mathrm{tr}}-\Phi w_{\mathrm{MAP}}\bigr\|_{2}^{2},$
$w_{\mathrm{MAP}}$ $=\bigl(\Phi^{\top}\Phi+\lambda I\bigr)^{-1}\Phi^{\top}y_{\mathrm{tr}}.$
For numerical stability and efficiency, we avoid forming $\Sigma$ explicitly. Let
$M\;=\;\lambda I\;+\;\tfrac{1}{\sigma^{2}}\,\Phi^{\top}\Phi\;=\;LL^{\top}$
be the Cholesky factorization. Then the epistemic term is evaluated without materializing $\Sigma$ as
$\phi(x)^{\top}\Sigma\,\phi(x)\;=\;\bigl\|L^{-1}\phi(x)\bigr\|_{2}^{2},$
which we compute with two triangular solves.
@@ 3.3 Two-Sided Posterior CDF Score
where $F_{\text{post}}$ is the CDF of $\mathcal{N}(\hat{\mu}(x),\hat{v}(x))$. Intuitively, $s$ is large for outcomes typical under the predictive distribution (near its center) and small for tail outcomes; it thus aligns nonconformity with the full posterior shape instead of only point residuals. For the Gaussian case, if $z=\frac{y-\hat{\mu}(x)}{\sqrt{\hat{v}(x)}}$ and $\Phi(\cdot)$ is the standard normal CDF, then $s(x,y)=\min\{\Phi(z),1-\Phi(z)\}$.
@@ 3.4 Split-Conformal Calibration
Compute scores on the calibration set $S_{\text{cal}}=\{s(x_{i},y_{i})\}_{(x_{i},y_{i})\in\mathcal{D}_{\text{cal}}}$ and take the rank-based threshold
$t\;=\;\text{Quantile}_{\alpha}\big(S_{\text{cal}}\big),$
where $\alpha=\frac{\lceil(|\mathcal{D}_{\text{cal}}|+1)(1-\text{target\_cov})\rceil}{|\mathcal{D}_{\text{cal}}|+1}$ (standard split-conformal choice). For a new $x$, the conformal prediction set is
$\mathcal{C}(x)\;=\;\big\{y:\;s(x,y)\geq t\big\}.$
With Gaussian $F_{\text{post}}$, $\mathcal{C}(x)$ is a central interval:
$\mathcal{C}(x)=\Big[\hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(t),\ \hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(1-t)\Big].$
@@ 3.5 CLAPS Procedure
A compact posterior-aware conformal routine: fit an LLLA posterior on the head, score calibration samples by two-sided CDF centrality, take the conformal quantile, and output central intervals. See Algorithm 1 for a summary.
@@
8:for all test $x$ do
9:  $\mu\leftarrow\phi(x)^{\top}w_{\mathrm{MAP}}$, $v\leftarrow\sigma^{2}+\|L^{-1}\phi(x)\|_{2}^{2}$
10:  $\mathcal{C}(x)\leftarrow\big[\mu+\sqrt{v}\,\Phi^{-1}(t),\ \mu+\sqrt{v}\,\Phi^{-1}(1-t)\big]$
11:end for
@@ 3.6 Complexity and Deployment
If $d$ is the last-layer width and $n$ the training size, forming $M$ costs $O(nd^{2})$ and its Cholesky $O(d^{3})$ once per model. Per-query costs are $O(d^{2})$ for $\hat{\mu}(x)$ and $O(d^{2})$ (or $O(d)$ with cached $L^{-1}$-vector solves) for $\hat{v}(x)$. CLAPS adds negligible overhead to standard same-backbone pipelines while providing a posterior-aware score amenable to fast batch inference and plug-and-play conformal calibration.
@@ 4 Theory
Let $w\mid\mathcal{D}_{\mathrm{tr}}\sim\mathcal{N}(w_{\mathrm{MAP}},\Sigma)$ with $\Sigma=(\lambda I+\sigma^{-2}\Phi^{\top}\Phi)^{-1}$ and $\Phi=[\phi(x)]_{(x,y)\in\mathcal{D}_{\mathrm{tr}}}$. The posterior predictive is
@@
and numerically $\phi(x)^{\top}\Sigma\phi(x)=\lVert L^{-1}\phi(x)\rVert_{2}^{2}$ for $LL^{\top}=\lambda I+\sigma^{-2}\Phi^{\top}\Phi$.
@@
For the Gaussian case, with $z=(y-\hat{\mu}(x))/\sqrt{\hat{v}(x)}$, we have $s(x,y)=\min\{\Phi(z),1-\Phi(z)\}$. Let $t$ be the split-conformal rank quantile of $\{s(x_{i},y_{i})\}_{(x_{i},y_{i})\in\mathcal{D}_{\mathrm{cal}}}$ at level $\alpha$.
@@
Under Assumption 1, the CLAPS interval
$\mathcal{C}(x)=\Big[\hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(t),\;\hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(1-t)\Big]$
satisfies $\Pr\{\,y\in\mathcal{C}(x)\,\}\geq 1-\alpha$ marginally over the calibration set and the test point.
@@
For any fixed $x$ with Gaussian predictive, the superlevel set $\{y:s(x,y)\geq t\}$ is the shortest (HPD) interval of posterior mass 12t1-2t. Consequently, CLAPS aligns the conformal acceptance region with the posterior notion of “typicality” rather than raw residual magnitude.
@@
Fix $x$ and suppose $F_{\mathrm{post}}(\cdot\mid x)$ equals the true conditional distribution. Among scores whose superlevel sets are nested in $t$, CLAPS yields intervals of minimal Lebesgue measure $|\mathcal{C}(x)|$ subject to $\Pr\{y\in\mathcal{C}(x)\mid x\}\geq 1-\alpha$.
@@
Shortest-probability sets at a fixed mass are superlevel sets of the density (HPD). For symmetric unimodal Gaussians, HPD coincides with central quantile intervals, which are induced by thresholding the two-sided CDF. Conformal calibration transfers the oracle probability constraint to finite-sample marginal coverage while preserving the nesting, hence minimal expected width under correct specification. ∎
@@
As $n=\lvert\mathcal{D}_{\mathrm{tr}}\rvert\to\infty$ with fixed backbone width, the spectrum of $\Phi^{\top}\Phi$ grows and $\Sigma=(\lambda I+\sigma^{-2}\Phi^{\top}\Phi)^{-1}\to 0$. Hence $\phi(x)^{\top}\Sigma\phi(x)\to 0$ for all $x$ and $\hat{v}(x)\to\sigma^{2}$ (last-layer epistemic collapse). In this regime CLAPS reduces to conformalization of a homoscedastic Gaussian predictor, and its width advantage over residual-based scores diminishes.
@@
If $r$ and $\mathrm{trace}(\Sigma)$ are nontrivial while $\rho$ is modest, prefer CLAPS; if $r\approx 0$ (contraction) but $\rho$ is substantial (heteroscedasticity), prefer scale-learning baselines (e.g., normalized-CP/CQR). This rule is consistent with Theorems 4.7 and 4.9.

### v3

@@ 2 Related Work
Split conformal regression calibrates residual-shaped scores to guarantee marginal coverage, and fold-reuse variants such as jackknife+ and cross-fitting improve data efficiency (Barber et al., 2021; Angelopoulos et al., 2023). Recent work emphasizes efficiency—minimizing expected length or controlling a user-chosen risk—while preserving guarantees, including optimization-driven procedures and general risk-control frameworks (Kiyani et al., 2024; Angelopoulos et al., 2022; Overman et al., 2024). Concurrently, methods move beyond absolute residuals toward distribution-shaped scores that leverage estimates of P(Y|X)P(Y\,|\,X) (e.g., CDF-/quantile-based and probabilistic variants) or reframe regression as classification to better capture heteroscedasticity, skewness, and multimodality (Chernozhukov et al., 2021; Romano et al., 2019; Guha et al., 2024; Plassier et al., 2024a). We adopt the same rank-based split calibration rule but define the conformity via a symmetric CDF from a predictive distribution (instantiated later with LLLA), placing our approach on the distribution-shaped branch; for completeness we also compare to interval-thresholding baselines tailored for width efficiency (Luo and Zhou, 2025).
@@
A major route to width efficiency under input-dependent noise is to learn scale or quantiles and calibrate them with conformal ranks. Conformalized Quantile Regression (CQR) replaces absolute-residual scores by a max-quantile deviation, achieving robustness to heteroscedasticity and skewed tails (Romano et al., 2019). Building on this idea, boosted or reweighted schemes further reduce interval length by improving the base quantile fits while preserving split-calibrated guarantees (Xie et al., 2024). Localization and weighting adapt calibration to covariate neighborhoods or strata, yielding relaxed conditional guarantees and improved coverage balance across subpopulations (Hore and Barber, 2023; Bhattacharyya and Barber, 2024). When conditional distributions are multi-modal/asymmetric, reframing regression targets (e.g., via discretization or surrogate tasks) can enhance quantile learning before conformalization (Guha et al., 2024). Recent general treatments study heterogeneity and non-exchangeability directly, proposing efficient conformal procedures with validity under drift and data heterogeneity (Plassier et al., 2024b). In survival/structured settings, conformalized quantile ideas are extended with monotonicity-aware or task-specific heads to stabilize learned quantiles while keeping rank calibration (Qi et al., 2024). We adopt CQR and localized/weighted variants as primary baselines; our method differs by deriving a posterior-aware, CDF-symmetric score, whereas these approaches learn input-dependent scales/quantiles and then apply the same split rank rule.
@@
Posterior-aware conformal methods use a predictive distribution to define CDF/HPD-style conformity scores, then retain split/rank calibration. Recent advances show how modelling the conditional distribution improves conditional validity and efficiency, e.g., clustering- or distribution–driven calibration (Zhang and Candès, 2024; Plassier et al., 2024a; Chernozhukov et al., 2021). In neural networks, a lightweight route to such predictive distributions is to endow only the final linear layer with a Bayesian treatment. Last-layer Laplace and its modern variants provide Gaussian posterior predictives around the MAP solution with negligible retraining cost (Kristiadi et al., 2020; Daxberger et al., 2021). Beyond the classical diagonal/empirical-Hessian forms, recent work improves the fidelity–cost trade-off via variational last layers, function-space priors, and geometry-aware approximations (Harrison et al., 2024; Cinquin et al., 2024; Bergamin et al., 2023). We adopt this last-layer Bayesianization to instantiate a symmetric CDF score from the LLLA posterior predictive on a shared MLP backbone, then calibrate by ranks; this places our method at the intersection of posterior-aware conformal scoring and efficient Laplace-based uncertainty, and is complementary to non-Bayesian distributional surrogates (Zhang and Candès, 2024; Plassier et al., 2024a).
@@
Beyond score design, three complementary axes improve efficiency and coverage balance. First, localization tailors calibration to neighborhoods of the test covariate, yielding relaxed conditional guarantees and better subgroup balance (Guan, 2023; Hore and Barber, 2023). Second, weighting focuses coverage where distribution shift or stratification matters—either by importance-style reweighting or group-aware corrections—without changing the base score (Bhattacharyya and Barber, 2024). Third, bootstrap ensembles aggregate diverse learners or resamples to stabilize interval length under model misspecification or dependence, including sequential/time-series settings (Xu and Xie, 2021, 2023; Gupta et al., 2022; Qian et al., 2024; Rivera et al., 2024). These lines are orthogonal to our posterior-aware scoring: our method can plug into localized or weighted calibration, and can be paired with ensemble wrappers to trade off width and robustness while preserving split/rank guarantees.
@@ 3 Method
We propose CLAPS, a posterior-aware conformal regression method that combines a Last-Layer Laplace Approximation (LLLA) with split-conformal calibration. CLAPS aligns the conformity score with the model’s posterior predictive shape and then uses rank-based calibration to guarantee finite-sample coverage.
@@ 3.2 Last-Layer Laplace Approximation (LLLA)
We place a Gaussian prior $w\sim\mathcal{N}(0,\lambda^{-1}I)$ and assume homoscedastic Gaussian noise $\tilde{y}\mid x,w\sim\mathcal{N}(\phi(x)^{\top}w,\sigma^{2})$, where $\tilde{y}$ denotes the standardized targets on the training set. Let $\mathbf{H}\in\mathbb{R}^{n\times d}$ be the design matrix with rows $\phi(x_{i})^{\top}$ for $(x_{i},y_{i})\in\mathcal{D}_{\mathrm{tr}}$, and let $\tilde{y}\in\mathbb{R}^{n}$ be the corresponding standardized targets.
Implementation (ridge mean + plug-in variance).
In practice we compute the last-layer mean via ridge regression on standardized targets:
$w_{\mathrm{MAP}}=\arg\min_{w}\frac{1}{2}\|\tilde{y}-\mathbf{H}w\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2}=(\mathbf{H}^{\top}\mathbf{H}+\lambda I)^{-1}\mathbf{H}^{\top}\tilde{y}$.
We estimate the noise variance by the training residual MSE,
$\hat{\sigma}^{2}=\frac{1}{n}\,\|\tilde{y}-\mathbf{H}w_{\mathrm{MAP}}\|_{2}^{2},$
and form the Gaussian posterior covariance (plug-in):
$\Sigma=\Bigl(\lambda I+\tfrac{1}{\hat{\sigma}^{2}}\mathbf{H}^{\top}\mathbf{H}\Bigr)^{-1}.$
This yields a Gaussian posterior predictive on the standardized scale,
$\tilde{y}\mid x,\mathcal{D}_{\mathrm{tr}}$ $\sim\mathcal{N}\!\bigl(\hat{\mu}(x),\,\hat{v}(x)\bigr),$
$\hat{\mu}(x)$ $=\phi(x)^{\top}w_{\mathrm{MAP}},$
$\hat{v}(x)$ $=\hat{\sigma}^{2}+\phi(x)^{\top}\Sigma\,\phi(x).$
Bayesian interpretation.
Under the Gaussian prior and likelihood above, the posterior over $w$ is Gaussian with mean $w_{\mathrm{MAP}}$ and covariance $\Sigma$. (The expression uses a plug-in estimate $\hat{\sigma}^{2}$ for numerical simplicity.) Unless otherwise stated we set the prior precision to $\lambda=1.0$.
Efficient evaluation.
For numerical stability and efficiency we avoid forming $\Sigma$ explicitly. Let
$M=\lambda I+\tfrac{1}{\hat{\sigma}^{2}}\mathbf{H}^{\top}\mathbf{H}=LL^{\top}$
be the Cholesky factorization. Then the epistemic term is computed as
$\phi(x)^{\top}\Sigma\,\phi(x)=\|L^{-1}\phi(x)\|_{2}^{2},$
using two triangular solves.
Notation.
For notational simplicity, in the remainder of the paper we write $y$ for the standardized targets $y_{\mathrm{tr}}$ (or $\tilde{y}$), and write $\sigma^{2}$ for the plug-in estimate $\hat{\sigma}^{2}$ computed from training residuals in this section.
@@ 3.3 Two-Sided Posterior CDF Score
where $F_{\text{post}}$ is the CDF of $\mathcal{N}(\hat{\mu}(x),\hat{v}(x))$. Intuitively, $s$ is large for outcomes typical under the predictive distribution (near its center) and small for tail outcomes. For the Gaussian case, if $z=\frac{y-\hat{\mu}(x)}{\sqrt{\hat{v}(x)}}$ and $\Phi(\cdot)$ is the standard normal CDF, then $s(x,y)=\min\{\Phi(z),1-\Phi(z)\}$.
@@ 3.4 Split-Conformal Calibration
Let $m=|\mathcal{D}_{\mathrm{cal}}|$ and compute calibration scores $S_{\text{cal}}=\{s(x_{i},y_{i})\}_{(x_{i},y_{i})\in\mathcal{D}_{\text{cal}}}$. Sort them as $s_{(1)}\leq\cdots\leq s_{(m)}$. Given a nominal miscoverage level $\alpha\in(0,1)$ (so the target coverage is $1-\alpha$), define the rank index
$k\;=\;\left\lceil(m+1)\alpha\right\rceil,\qquad t\;=\;s_{(k)}.$
Because $s$ is a conformity score (larger values indicate more typical outcomes), we calibrate using the lower $\alpha$-quantile so that $\Pr\{s(x_{\ast},y_{\ast})\geq t\}\geq 1-\alpha$. For a new $x$, the conformal prediction set is
$\mathcal{C}_{\alpha}(x)\;=\;\big\{y:\;s(x,y)\geq t\big\}.$
With Gaussian $F_{\text{post}}$, $\mathcal{C}_{\alpha}(x)$ is a central interval:
$\mathcal{C}_{\alpha}(x)=\Big[\hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(t),\ \hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(1-t)\Big].$
@@ 3.5 CLAPS Procedure
A compact posterior-aware conformal routine: fit an LLLA posterior on the head, score calibration samples by two-sided CDF centrality, take the conformal rank threshold, and output central intervals. See Algorithm 1 for a summary.
@@
7: Sort $\{s_{i}\}_{\mathcal{D}_{\mathrm{cal}}}$ into $s_{(1)}\leq\cdots\leq s_{(m)}$ where $m=|\mathcal{D}_{\mathrm{cal}}|$
@@
9:for all test $x$ do
10:  $\mu\leftarrow\phi(x)^{\top}w_{\mathrm{MAP}}$, $v\leftarrow\sigma^{2}+\|L^{-1}\phi(x)\|_{2}^{2}$
11:  $\mathcal{C}_{\alpha}(x)\leftarrow\big[\mu+\sqrt{v}\,\Phi^{-1}(t),\ \mu+\sqrt{v}\,\Phi^{-1}(1-t)\big]$
12:end for
@@ 3.6 Complexity and Deployment
If $d$ is the last-layer width and $n$ the training size, forming $\mathbf{H}^{\top}\mathbf{H}$ costs $O(nd^{2})$ and its Cholesky (via $M$) costs $O(d^{3})$ once per model. Per-query costs are $O(d)$ for $\hat{\mu}(x)=\phi(x)^{\top}w_{\mathrm{MAP}}$ and $O(d^{2})$ for $\hat{v}(x)$ (two triangular solves to compute $\|L^{-1}\phi(x)\|_{2}^{2}$). CLAPS adds negligible overhead to standard same-backbone pipelines while providing a posterior-aware score amenable to fast batch inference and plug-and-play conformal calibration.
@@ 4 Theory
Let $w\mid\mathcal{D}_{\mathrm{tr}}\sim\mathcal{N}(w_{\mathrm{MAP}},\Sigma)$ with $\Sigma=(\lambda I+\sigma^{-2}\mathbf{H}^{\top}\mathbf{H})^{-1}$ and $\mathbf{H}=[\phi(x)^{\top}]_{(x,y)\in\mathcal{D}_{\mathrm{tr}}}$. The posterior predictive is
@@
and numerically $\phi(x)^{\top}\Sigma\phi(x)=\lVert L^{-1}\phi(x)\rVert_{2}^{2}$ for $LL^{\top}=\lambda I+\sigma^{-2}\mathbf{H}^{\top}\mathbf{H}$.
@@
For the Gaussian case, with $z=(y-\hat{\mu}(x))/\sqrt{\hat{v}(x)}$, we have $s(x,y)=\min\{\Phi(z),1-\Phi(z)\}$, where $\Phi(\cdot)$ denotes the standard normal CDF. Let $t$ be the split-conformal rank quantile of $\{s(x_{i},y_{i})\}_{(x_{i},y_{i})\in\mathcal{D}_{\mathrm{cal}}}$ at level $\alpha$.
@@
Under Assumption 4.3, let $m=|\mathcal{D}_{\mathrm{cal}}|$ and let $t$ be the split-conformal rank threshold computed from $\mathcal{D}_{\mathrm{cal}}$ at level $\alpha$. Then the CLAPS interval
$\mathcal{C}_{\alpha}(x)=\Big[\hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(t),\;\hat{\mu}(x)+\sqrt{\hat{v}(x)}\,\Phi^{-1}(1-t)\Big]$
satisfies finite-sample marginal coverage conditional on the training data:
$\Pr\!\bigl\{\,y_{\ast}\in\mathcal{C}_{\alpha}(x_{\ast})\ \big|\ \mathcal{D}_{\mathrm{tr}}\,\bigr\}\ \geq\ 1-\alpha,$
where the probability is over the exchangeable draws of $\mathcal{D}_{\mathrm{cal}}$ and $(x_{\ast},y_{\ast})$ given $\mathcal{D}_{\mathrm{tr}}$. In particular, $\Pr\{\,y_{\ast}\in\mathcal{C}_{\alpha}(x_{\ast})\,\}\geq 1-\alpha$ marginally.
@@
For any fixed $x$ with Gaussian predictive, the superlevel set $\{y:s(x,y)\geq t\}$ is the shortest (HPD) interval of posterior mass 12t1-2t. Consequently, CLAPS aligns the conformal acceptance region with the posterior notion of ”typicality” rather than raw residual magnitude.
@@
Fix $x$ and suppose $F_{\mathrm{post}}(\cdot\mid x)$ equals the true conditional distribution of $Y\mid X=x$. In our setting $F_{\mathrm{post}}(\cdot\mid x)$ is Gaussian (from the LLLA predictive), so this oracle condition means the true conditional law $Y\mid X=x$ is Gaussian with mean $\hat{\mu}(x)$ and variance $\hat{v}(x)$. Among score-based prediction sets $\mathcal{C}(x)=\{y:s(x,y)\geq t\}$ whose acceptance regions are nested in $t$, CLAPS yields intervals of minimal Lebesgue measure $|\mathcal{C}(x)|$ subject to $\Pr\{Y\in\mathcal{C}(x)\mid X=x\}\geq 1-\alpha$.
@@
Under correct specification, the shortest conditional $(1-\alpha)$ prediction set at a fixed $x$ is the highest-posterior-density (HPD) set, i.e., a superlevel set of the conditional density. For symmetric unimodal Gaussians, the HPD set coincides with the central quantile interval, which is equivalently obtained by thresholding the two-sided posterior CDF score. Split-conformal rank calibration then chooses a data-dependent threshold $t$ to achieve finite-sample marginal coverage $1-\alpha$ without changing the nested superlevel-set form, so the resulting CLAPS sets inherit oracle efficiency within this nested family under correct specification. ∎
Remark 4.8 (Nested acceptance sets).
Nestedness means that for $t_{1}\geq t_{2}$, $\{y:s(x,y)\geq t_{1}\}\subseteq\{y:s(x,y)\geq t_{2}\}$, i.e., raising the threshold only shrinks the accepted set. Theorem 4.7 states efficiency within this nested family; without nestedness, set-length optimality need not hold.
@@
Let $\mathbf{H}=[\phi(x_{i})^{\top}]_{i=1}^{n}\in\mathbb{R}^{n\times d}$ be the last-layer design matrix built from $\mathcal{D}_{\mathrm{tr}}$, and assume $\Gamma:=\mathbb{E}[\phi(x)\phi(x)^{\top}]$ exists with $\lambda_{\min}(\Gamma)>0$. As $n=\lvert\mathcal{D}_{\mathrm{tr}}\rvert\to\infty$ with fixed backbone width $d$, the spectrum of $\mathbf{H}^{\top}\mathbf{H}$ grows and $\Sigma=(\lambda I+\sigma^{-2}\mathbf{H}^{\top}\mathbf{H})^{-1}\to 0$. Hence $\phi(x)^{\top}\Sigma\phi(x)\to 0$ for all $x$ and $\hat{v}(x)\to\sigma^{2}$ (last-layer epistemic collapse). In this regime CLAPS reduces to conformalization of a homoscedastic Gaussian predictor, and its width advantage over residual-based scores diminishes.
@@
If $r$ and $\mathrm{trace}(\Sigma)$ are nontrivial while $\rho$ is modest, prefer CLAPS; if $r\approx 0$ (contraction) but $\rho$ is substantial (heteroscedasticity), prefer scale-learning baselines (e.g., normalized-CP/CQR). This rule is consistent with Theorems 4.7 and 4.10.